# Learn, Check, Test – Security Testing Using Automata Learning and Model Checking

### v1

@@ 5.4 Illustrative Example
Environment(){
self.req();
msgsrv req() {
omega2set=false;
int data = (0);
switch(data) {
case 0: system.sigma1(); break;
msgsrv omega1(){
self.req();
msgsrv omega2(){
omega2set=true;
self.req();
@@
msgsrv sigma1(){
@@
state=1;
p=true;
environment.omega1();
@@ 5.5 Altering the Model
Since in Rebeca a model’s behavior is defined similar to a programming language, altering it is trivial. Furthermore, Rebeca possess an operator for non-determinism (the ?? operator). On the one hand, we can manually add any arbitrary fault that might be interesting to investigate and examine its impact on keeping the properties. On the other hand, this allows for re-introducing some non-deterministic behavior that is abstracted away during the learning process (see Section 2.2.1[52]. By altering the template mentioned in Section 5.3, we just introduce a timeout server on the sender side and alter each input message server on the receiver to trigger a timeout with a certain probability [^14] Realized with a random (??) pick of an integer. If no timeout occurs, the logic continues. Otherwise, we reset the state to initial and call the timeout function on the sender, which sets a temporary proposition for timeout, to allow for checking for it. Since we can set the timeout probability to an arbitrary value, we can play with different scenarios to be checked. We also can convert the resulting Rebeca code back to a Mealy machine, so we possess a model that can contain timeouts and errors, just as the originally learned model would have [^15] If the resulting Rebeca code contains non-determinism, the resulting model is not technically a Mealy machine, but a Mealy-styled non-deterministic state model, for it defines multiple transitions for a given input per state to represent the different possibilities.. Formally, we alter the Rebeca template by extending $M_{env}$ with a timeout method, that sets a state variable ${timeout}\in V_{env}$ (to perform checks for timeouts if needed) and calls the request method. Further, we alter each method $m_{sys}\in M_{sys}$ to randomly (with a settable parameter determining the probability) reset the state identifier $q$ to the initial state $q_{1}$ and call the timeout method ($\langle r_{sys},r_{env},{timeout}\rangle$), instead of the output method ($M_{env}(\omega)$ derived from $\lambda(q,\sigma)$.
@@ 5.6 Verifiying the Code
In order to assure the Rebeca code is faithful to the learned model, we check their equivalence. We generate the full state space of the Rebeca model. For that, we run the model checker without any properties to check, so, the model checker does not stop at a violation. This way the full state space is generated, which we export to a Linear Transition System (LTS) representation in the GraphViz format. Based on a previous work in [53], we then use an algorithm that collapses the state space LTS into a Mealy machine and check the bisimilarity of the converted Rebeca Mealy machine with the originally learned model. The Rebeca model checking tool (RMC) supports Linear Temporal Logic (LTL) to formulate properties. For security checking, these LTL form should describe security attributes.
@@ 6.1 Defining Generic Properties
Since the requirements in Section 3 are very generic, we can also define generic properties to check them. The become more specific in conjunction with the CPMs (Section 5.1), which define the meaning of propositions used in the property in the context of a specific protocol [^16] If a proposition used in a property is not defined in the respective CPM, we define it as false($\bot$) for the respective property.. For the sake of readability, we restrict our definitions of access to read operations but write, update, and similar operations can be defined the same way.
@@ 6.1.1 Authentication
For Requirement R3.1, we define an authenticated state as a state with the proposition authenticated set to false, a when an access operation on a protected resource occurs, then the operation must not return a positive response (access ok).
@@
Here, we define AUTH{AUTH} for authenticated, PROT{PROT} for protected, and ACCESSOK{ACCESSOK} for a positive response [^18] Example for non-existing protocols: if we have a protocol with no protected resources, we set $PROT:=\bot$, which means that reading is always allowed according to Property P6.1.1. (i.e., a successful read operation).
@@ 6.1.2 Confidentiality
For checking the confidentiality (Requirement R3.2), we assume the presence of unsecured and secured operations. For the access operation as used in Section 6.1.1, we therefore define two exemplary subsets: an unsecured read $UREAD$ and a secured read $SREAD$. Since $READ$ is a superset, in the respective context maps we then define that $READ$ will always be set along with $UREAD$ or $SREAD$ is set (using the same conditions, i.e., $c_{READ}\leftrightarrow c_{UREAD}\lor c_{SREAD}$). Given requirement R3.2, we define that an $UREAD$ may not succeed when $PROT$ is true. Please note that we fail to apply this property to the UDS case, since there is no encrypted communication on the CAN bus that servers as a medium for the UDS protocol [^19] While there are some proprietary and research solutions for encrypted CAN, none of them is widely adopted (i.e., used by a larger number of major OEMs). An industry standard gaining momentum, the AUTOSAR Secure On-board Communication (SecOC) [54], only provides authentication and integrity, but not confidentiality..
@@ 6.1.3 Privilege Levels
We assume that there are areas that go beyond a basic protection level and that there additional protection measures (e.g., different type or level of authentication) have to be taken (Requirement R3.3). We define this higher protection level as critical resources ($CRIT$) and the additional authentication as privileged ($PRIV$). We further assume that since it is a higher security level privileged authentication ($PRIV$) also includes normal authentication ($AUTH$).
@@ 6.1.4 Key Validity
A property for using a valid key (Requirement R3.4) is straightforward, since it only has to define that an invalid key must not be accepted in the entire system.
@@ 6.2 Checking the Properties
This section defines and explains the CPMs for eMRTDs and UDS, along with defining some protocol-specific properties that allow for checking the protocol flow.
6.2.1 eMRTDs
@@
For eMRTDs, we define $PROT$ as $DF$ input in $C_{g}$ of the CPM for eMRTDs (Table 1), since the ICAO standard defines the applications to be in a protected zone [^20] Note: we only set DF for selecting LDS1, which corresponds to the eMRTD application. Technically, DF would also be set when selecting other applications (i.e., LDS2). However, we then would design the CPM to distinguish between DFs. This is not necessary in our case, since we only use one., while $AUTH$ is gained by performing a basic authentication (BAC) operation. As $ACCESSOK$ we define a successfully performed file selection in $C_{\tau}$. This means that for P6.1.1 a BAC must be performed before selecting any file inside an application.
@@
For $UREADOK$, we define an unsecured read binary operation. This means that for P6.1.2, only secured read binary operations are allowed.
@@
We define $CRIT$ as a successful selection of data groups 2 or 3 in $C_{\tau}$. These contain biometric data, which is why the ICAO standard requires and additional (particularly terminal) authentication. Since, we do not have an additional authentication method implemented, $PRIV$ is set to false ($\bot$). This means that P6.1.3 only holds when these sensitive data groups cannot be selected in the complete model.
@@
Lastly, we define $INVKEYOK$ for P6.1.4 as any successful secured operation that has been carried out using an old key or an all-zero key in $C_{\tau}$. We realize these by dedicated input symbols (see Section 4.1.1).
@@
We also use a couple of complimentary properties for checking eMRTDs. SecureRead: $\square(\neg(SREADOK\land\neg(DF\land AUTH\land EF)))$, which reads as a secure read operation (SREADOK) can only be successful if authenticated and a protected resource is selected ($DF\land EF$). This also creates the need for a normal read operation not to work on protected resources: PlainRead: $\square(\neg(UREADOK\land(\neg EF\lor DF)))$, which reads as a successful plain read (READOK) cannot happen without a selected resource outside of LDS1 ($\neg EF$$\lor DF$). We can further specify a secure select process with the following property: SecureReadFollowsSecureSelect: $((\neg SREADOK)\ \mathcal{U}\$$\text{SSELEFOK})\lor\square(\neg SREADOK)$. Table 1 gives an overview of the CPM.
6.2.2 UDS
@@
Authentication in UDS works over SecurityAccess and subsequent SecurityAccessWithKey. We therfore define the $AUTH$ proposition after the second step has been successfully exectued. As a canary for P6.1.1, we perform a reading operation on a protected resource. We choose to start the Check Programming Dependencies (0xFF01) routine. This routine does not alter the firmware and is standardized [8]. Furthermore, it is industry practice that this route is protected by a Security Access routine [55]. Therefore we set both $PROT$ and $ACCESSOK$ as a $c_{\tau}$ when starting this routine. For P6.1.2, set perform a direct memory read operation to set $UREADOK$. Unfortunately, we don not have the knowledge of which memory are actually protected (as a workaround defining the CheckASWBit routine could also be defined as $PROT$). For P6.1.3, we define the RequestDownload routine as $CRIT$, as it allows to actually flash the device, while $PRIV$ his defined as SecurityAcccess with a higher level. Lastly for P6.1.4, $INVKEYOK$ is defined as SecurityAcccess with a wrong key (as an own input symbol). Table 2 shows the CPM for UDS.
@@ 7 Testing
With a property violation discovered by the learning and checking procedure, the only thing left is to verify the finding on the actual system. Due to relying on conformance testing for equivalence queries, the automata learning process maintains a residual risk for an inaccurate model. We therefore test the trace of a found violation with the life system to exclude false positives. If the actual system behavior matches the one predicted by the model, the violation is confirmed. Otherwise, the trace can be fed back to the learning system to refine the model. The verification of found traces is trivial, as the same tool set from the learning can be used for concretizing the input and interfacing with the SUL. Therefore, this combination can be seen as method for test case generation.
@@ 8 Evaluation
To show the practical usability and versatility of the described process, we evaluate it on two devices (an eMRTD and an automotive control unit), each speaking one of the communication protocols described earlier in this paper (NFC and UDS). Though the use cases are very different, in both cases the model checking is very similar. At the present abstraction level, we check generic security requirements (Section 3) using according LTL properties (Section 6.1).
@@ 8.1 Electronically Machine-Readable Travel Document
For the eMRTD use case, we scrutinize two different Austrian passports, on elder, expired and one current (see a process based on equivalence checking in [34]). The model was learned via LearnLib using the TTT algorithm (with binary closure) and a minimum input trace length of 40 symbols, a maximum of 50, and 150 random walk conformance tests as equivalence oracle.

### v2

@@ 6.3 Rebeca Code Transformation
ENVIRONMENT() { self.start(); }
void start() { system.req(); }
msgsrv err(){ system.req(); }
msgsrv ok(){ system.req(); }
@@
statevars {
boolean prot, auth;
int state;
boolean error, accessok;
SYSTEM() {
prot=false;
auth=false;
state=0;
msgsrv req() {
error=false;
accessok=false;
int data =?(0,1);
switch(data) {
case 0: self.cred(); break;
case 1: self.read(); break;
default: self.ERR();
msgsrv ERR(){
error = true;
msgsrv read(){
@@
state=0;
environment.err();
@@
state=1;
accessok=true;
environment.ok();
msgsrv cred(){
if(state==0) {
auth=true;
state=1;
environment.ok();
} else
if(state==1) {
auth=false;
@@ 6.4 Altering the Model
Since in Rebeca a model’s behavior is defined similar to a programming language, altering it is trivial. Furthermore, Rebeca possess an operator for non-determinism (the ??-operator). On the one hand, we can manually add any arbitrary fault that might be interesting to investigate and examine its impact on keeping the properties. On the other hand, this allows for re-introducing some non-deterministic behavior that is abstracted away during the learning process [47]. For instance, we reintroduced timeouts that have been abstracted, by altering the template mentioned in Section 6.1. We just introduce a timeout server on the sender side and alter each input message server on the receiver to trigger a timeout with a certain probability, realized with a random (??) pick of an integer. If no timeout occurs, the logic continues. Otherwise, we reset the state to initial and call the timeout function on the sender, which sets a implicit state proposition for timeout, to allow for checking for it. Since we can set the timeout probability to an arbitrary value, we can play with different scenarios to be checked. We also can convert the resulting Rebeca code back to a Mealy machine, so we possess a model that can contain timeouts and errors, just as the originally learned model would have. If the resulting Rebeca code contains non-determinism, the resulting model is not technically a Mealy machine, but a Mealy-styled non-deterministic state model, for it defines multiple transitions for a given input per state to represent the different possibilities. Formally, we alter the Rebeca template by extending $M_{env}$ with a timeout method, that sets a state variable ${timeout}\in V_{env}$ (to perform checks for timeouts if needed) and calls the request method. Further, we alter each method $m_{sys}\in M_{sys}$ to randomly (with a settable parameter determining the probability) reset the state identifier $q$ to the initial state $q_{1}$ and call the timeout method ($\langle r_{sys},r_{env},{timeout}\rangle$), instead of the output method ($M_{env}(\omega)$ derived from $\lambda(q,\sigma)$).
@@ 6.5 Verifiying the Code
In order to assure the Rebeca code is faithful to the learned model, we check their equivalence. We generate the full state space of the Rebeca model. For that, we run the model checker without any properties to check, so, the model checker does not stop at a violation. This way the full state space is generated, which we export to a Linear Transition System (LTS) representation in the GraphViz format. Based on a previous work in [30], we then use an algorithm that collapses the state space LTS into a Mealy machine and check the bisimilarity of the converted Rebeca Mealy machine with the originally learned model. This an extra optional step to provide assurance that no errors have been introduced in the transformation process.
@@ 7 Checking the Model
The Rebeca model checking tool (RMC) supports Linear Temporal Logic (LTL) to formulate properties. For security checking, these LTL formulas should describe security properties. Our approach allows us to use generic security properties, which we can concretize for specific state machines using CPMs. A property consists of propositions, which the CPM defines (using semantics from the CPM’s protocol context). This results in an AMM that includes the propositions, which the security properties we want to check consist of. In this chapter we begin with defining four basic, well-established security requirements (authentication, confidentiality, privilege levels, and key validity) and for each define a corresponding security property that we can check. Lastly, we briefly describe how to concretize the properties’ propositions using CPMs in two use cases (eMRTDs and UDS). Section 9 subsequently describes the practical application and evaluation in these two use cases.
7.1 Generic Security Requirements
In this section, we take some well established security properties [42] that are also aligned with our case studies’ standards documents (ICAO [15] and UDS [18]) and formulate them as checkable requirements in given-when-then [^2] Using given, when, and then as keywords in italics. format [5]  [^3] We use the terms MUST, MUST NOT, SHOULD, SHOULD NOT and MAY as defined by the Internet Engineering Task Force (IETF) [26], using the same CAPITALIZED style.. These requirements use attributes, we later (in Section 7.2) use as propositions [^4] Using small capitals style. for defining logic properties. This definition was performed manually, however, given the generality of the properties and their automated instantiation via CPMs we demonstrate below, the process can be considered a one-time effort. If other (either specific or generic) security properties are needed, we recommend to first state a semi-formalized requirement (e.g., in given-when-then format) analyze the components for attributable propositions and constitute the property to meet the requirement’s semantics. Subsequently, the properties can be instantiated by populating the propositions with a concrete system context with a CPM.
7.1.1 Authentication
To establish the authenticity as a property, we must require authentication of a user before accessing protected resources (not every resource must be protected, some may be openly accessible – this depends on the system’s security design). As a requirement, we define it as follows:
7.1.2 Confidentiality
Requirement R2:Given the system is in an arbitrary state, when an unsecured access operation on a protected resource occurs, then it SHOULD NOT return a positive response.
unsecured access thereby denotes a file operation without cryptographic security controls (i.e., encryption).
7.1.3 Privilege Levels
It is common practice for more complex systems and protocols to define more than one privilege level according to different roles and corresponding rights of an actor in order to provide access to more critical resources. This also applies for UDS, where this is implemented by different session and security levels [19]. Likewise, for eMRTDs, sensitve data have to be additionally protected via terminal authentication.
Requirement R3:Given the level of privilege is not sufficient, when an access operation on a critical resource occurs, then the operation MUST NOT be successful. A sufficient level of privileges also implies authentication.
7.1.4 Key Validity
Another important property of secure connections is key validity and recency (i.e., no newer key must exist). Since we do not consider a time component in the automata, we do not use recency in the sense of freshness. An invalid key in this sense is either a wrong key or an old key (not recent).
Requirement R4:Given any operation that requires a key, when an invalid key is provided, then the key MUST NOT be accepted.
@@ 7.2 Defining Generic Properties
Since the requirements in Section 7.1 are generic, we can also define generic properties to check them. Generally, we use the given and when parts as a premise and the then part as a consequence (except for the mapping from Requirement R7.1.4 to property P7.2.4, for which we trivially use no invalid keys as a property). These become more specific in conjunction with the CPMs, which define the meaning of propositions used in the property in the context of a specific protocol. If a proposition used in a property is not defined in the respective CPM, we have two options: a) leave it undefined, which creates an error in the model checking and makes the user aware of this fact or b) define it as false($\bot$) for the respective property. The latter option means that, for semantically sensible settings, our properties hold with undefined propositions, effectively ignoring the concerned proposition. For the sake of readability, we restrict our definitions of access to read operations; write, update, and similar operations can be defined the same way.
@@ 7.2.1 Authentication
For Requirement R7.1.1, we define an authenticated state as a state with the proposition authenticated set to true. If this is not the case, when an access operation on a protected resource occurs, then the operation must not return a positive response.
@@
Here, we define AUTH{AUTH} for authenticated, PROT{PROT} for protected, and ACCESSOK{ACCESSOK} for a positive response (i.e., a successful read operation). If we have a protocol with no protected resources, we can set $PROT:=\bot$, which means that reading is always allowed according to Property P7.2.1.
@@ 7.2.2 Confidentiality
For checking the confidentiality (Requirement R7.1.2), we assume the presence of unsecured and secured operations. For the access operation as used in Section 7.2.1, we therefore define two exemplary subsets: an unsecured read $UREAD$ and a secured read $SREAD$. Since $READ$ is a superset, in the respective context maps we then define that $READ$ will always be set along with $UREAD$ or $SREAD$ is set (using the same conditions, i.e., $c_{READ}=c_{UREAD}\lor c_{SREAD}$). Given requirement R7.1.2, we define that an $UREAD$ may not succeed when $PROT$ is true. Please note that we fail to apply this property to the UDS case, since there is no encrypted communication on the CAN bus that servers as a medium for the UDS protocol. While there are some proprietary and research solutions for encrypted CAN, none of them is widely adopted (i.e., used by a larger number of major OEMs). An industry standard gaining momentum, the AUTOSAR Secure On-board Communication (SecOC) [3], only provides authentication and integrity, but not confidentiality.
@@ 7.2.3 Privilege Levels
We assume that there are areas that go beyond a basic protection level and that there additional protection measures (e.g., different type or level of authentication) have to be taken (Requirement R7.1.3). We define this higher protection level as critical resources ($CRIT$) and the additional authentication as privileged ($PRIV$). We further assume that since it is a higher security level privileged authentication ($PRIV$) also includes normal authentication ($AUTH$).
@@ 7.2.4 Key Validity
A property for using a valid key (Requirement R7.1.4) is straightforward, since it only has to define that an invalid key must not be accepted in the entire system.
@@ 7.3 Checking the Properties
Once we have a Rebeca model according to a template (Definition 11) and respective properties, the Rebeca Model Checker (RMC) can check the model for these properties. Using a protocol-specific CPM, we define the propositions that constitute the properties in the context of a specific protocol. In our running example (see Figure 6), we defined the necessary propositions for Property P7.2.1 only (for the sake of simplicity), which holds for this example. If we alter the transition for $\sigma=READ$ in $q=s0$ from $\omega=ERR$ to $\omega=OK$, then respective implicit ($\tau$) state gets the proposition $ACCESSOK$, but not $AUTH$ and the property will be violated. Semantically the latter case would mean a successful access without authentication.
7.3.1 Checking eMRTDs
@@
For eMRTDs, we define $PROT$ as $DF$ input in $C_{g}$ of the CPM for eMRTDs (Table 3), since the ICAO standard defines the applications to be in a protected zone. We only set DF for selecting LDS1, which corresponds to the eMRTD application. Technically, DF would also be set when selecting other applications (i.e., LDS2). However, we then would design the CPM to distinguish between DFs. This is not necessary in our case, since we only use one. $AUTH$ is gained by performing a basic authentication (BAC) operation. As $ACCESSOK$ we define a successfully performed file selection in $C_{\tau}$. This means that for P7.2.1 a BAC must be performed before selecting any file inside an application.
@@
For $UREADOK$, we define an unsecured read binary operation. This means that for P7.2.2, only secured read binary operations are allowed.
@@
We define $CRIT$ as a successful selection of data groups 2 or 3 in $C_{\tau}$. These contain biometric data, which is why the ICAO standard requires and additional (particularly terminal) authentication. Since, we do not have an additional authentication method implemented, $PRIV$ is set to false ($\bot$). This means that P7.2.3 only holds when these sensitive data groups cannot be selected in the complete model.
@@
Lastly, we define $INVKEYOK$ for P7.2.4 as any successful secured operation that has been carried out using an old key or an all-zero key in $C_{\tau}$. We realize these by dedicated input symbols starting with WS (wrong key secured) and OS (old key secured).
@@
We also use a couple of complimentary properties for checking eMRTDs. SecureRead: $\square(\neg(SREADOK\land\neg(DF\land AUTH\land EF)))$, which reads as a secure read operation (SREADOK) can only be successful if authenticated and a protected resource is selected ($DF\land EF$). This also creates the need for a normal read operation not to work on protected resources: PlainRead: $\square(\neg(UREADOK\land(\neg EF\lor DF)))$, which reads as a successful plain read (READOK) cannot happen without a selected resource outside of LDS1 ($\neg EF$$\lor DF$). We can further specify a secure select process with the following property: SecureReadFollowsSecureSelect: $((\neg SREADOK)\ \mathcal{U}\ \allowbreak\text{SSELEFOK})\allowbreak\lor\square(\neg SREADOK)$. This means that a secured read operation may only follow after a secured selection operation has occurred, independent of the EF/DF properties in the AMM. The reason is to prevent errors like that the EF proposition has been obtained through an unsecured select EF. Table 3 gives an overview of the CPM.
7.3.2 Checking UDS
@@
Authentication in UDS works over SecurityAccess and subsequent SecurityAccessWithKey. We therefore define the $AUTH$ proposition after the second step has been successfully executed. To verify property P7.2.1, we perform a reading operation on a protected resource. We choose to start the Check Programming Dependencies (0xFF01) routine. This routine is standardized [19] and does not alter the firmware. Yet, it is industry practice that this route is protected by a Security Access routine as it is a preparatory action for flashing the system [56]. Therefore we set both $PROT$ and $ACCESSOK$ as a $c_{\tau}$ when starting this routine. For P7.2.2, set perform a direct memory read operation to set $UREADOK$. Unfortunately, we don not have the knowledge of which memory are actually protected (as a workaround defining the CheckASWBit routine could also be defined as $PROT$). For P7.2.3, we define the RequestDownload routine as $CRIT$, as it allows to actually flash the device, while $PRIV$ his defined as SecurityAcccess with a higher level. Lastly for P7.2.4, $INVKEYOK$ is defined as SecurityAcccess with a wrong key (as an own input symbol). Table 4 shows the CPM for UDS.
@@ 8 Testing
With a property violation discovered by the learning and checking procedure, the only thing left is to verify the finding on the actual system. Due to relying on conformance testing for equivalence queries, the automata learning process maintains a residual risk for an inaccurate model. We therefore test the trace of a found violation with the life system to exclude false positives. If the actual system behavior matches the one predicted by the model, the violation is confirmed. Otherwise, the trace can be fed back to the learning system to refine the model. We implement this step using a custom Oracle for the automata learner, where we can input the found counterexample for a new learning iteration. As a by-product, we use this custom oracle to improve the learning performance with a priori protocol knowledge by using test traces that resemble the standard behavior of a protocol first. The verification of found traces is trivial, as the same tool set from the learning can be used for concretizing the input and interfacing with the SUL. Therefore, this combination can be seen as method for test case generation.
@@ 9 Evaluation
To show the practical usability and versatility of the described process, we evaluate it on two devices (an eMRTD and an automotive control unit), each speaking one of the communication protocols described earlier in this paper (NFC and UDS). Though the use cases are very different, in both cases the model checking is very similar. At the present abstraction level, we check generic security requirements using according LTL properties. The properties and CPMs are pre-defined (whereby the specification is fairly easy), while the annotated Mealy is automatically generated (utilizing a CPM) from a fully automatically learned Mealy machine. Thus, we can claim that the complete process of learning and checking a real-life eMRTD and UDS-based system is largely automated, with very little human interaction necessary.
@@ 9.1 Electronically Machine-Readable Travel Document
For the eMRTD use case, we scrutinize two different Austrian passports, one elder, expired and one current (see a process based on equivalence checking in [29]). The model was learned via LearnLib using the TTT algorithm (with binary closure) and a minimum input trace length of 40 symbols, a maximum of 50, and 150 random walk conformance tests as equivalence oracle.